# Item Cluster-aware Prompt Learning for Session-based Recommendation

### v1

@@ 4.1.3. Evaluation Metrics
To evaluate the performance of the recommendation, we use two metrics widely adopted in SBR evaluation: Mean Reciprocal Rank (MRR@k) and Recall (Recall@k) following (Ruocco et al., 2017; Li et al., 2017; Quadrana et al., 2017), where $k=5$.
@@ 4.1.4. Implementation Details
We implement CLIP-SBR using a popular open-source recommender systems library RecBole [^4] https://recbole.io. All methods are optimized with the Adam optimizer, and we employ early stopping with a patience of 50 epochs to prevent overfitting, using MRR@5 as the indicator. The batch size is tuned from {64, 128, 256, 512, 1024} and learning rate from {0.001, 0.005, 0.01, 0.05}. For the Leiden algorithm used in item cluster detection, the resolution value is set to 1, with its impact discussed in detail in Section 4.2.4.
@@ 4.2.1. Overall Performance (RQ1)
Table 2 presents the experimental results of the eight baselines and their corresponding implementations using the CLIP-SBR framework on three real-world datasets. The best result in each column is highlighted in bold, and the top result within the intra-session category is underlined. We applied the CLIP-SBR framework to both intra-session and inter-session baselines to assess its impact. The results show that the CLIP-SBR leads to performance improvements across most of the baselines.
@@
In the intra-session baselines, we observed substantial improvements across the board, with the most significant performance gains in the NARM model. For the inter-session baselines, CLIP-SBR demonstrated notable improvements in GCSAN, while having a smaller impact on GCEGNN and no improvement on LESSR. Further analysis of the impact of applying CLIP-SBR to inter-session baselines is provided in Section 4.2.3.
@@
Moreover, the inter-session baselines (GCSAN, GCEGNN, LESSR) demonstrate comparable or even better performance than the intra-session baselines. Notably, GCEGNN outperforms all intra-session models on the Last.fm dataset, and LESSR achieves the highest metrics on the Xing dataset. These results highlight the potential of incorporating inter-session information to improve recommendation accuracy, emphasizing the importance of considering both intra-session and inter-session item relationships when developing robust session-based recommendation models.
@@
The application of the CLIP-SBR framework to the intra-session baselines resulted in performance improvements in all but one case. Specifically, NARM saw a substantial improvement of 45.30% on Last.fm, 20.11% on Xing, and 13.04% on Reddit, marking it as one of the most improved models. TAGNN also demonstrated significant gains, particularly on Last.fm with a 33.91% increase, 15.31% on Xing, and 2.33% on Reddit. Similarly, GRU4Rec achieved notable enhancements of 29.77% on Last.fm, 7.61% on Xing, and 11.10% on Reddit. CORE exhibited improvements of 1.92%, 16.65%, and 4.51% across the respective datasets, while SRGNN showed increases of 31.54%, 1.44%, and 4.26% on the three datasets. When comparing intra-session CLIP-SBR models with inter-session models, the CLIP-SBR models generally outperformed in many instances. Notably, CLIP-TAGNN consistently surpassed all inter-session models across. CLIP-SRGNN also outperformed inter-session models on Last.fm and Reddit dataset, and outperformed GCSAN in MRR@5 and Recall@5, and GCEGNN in MRR@5. This indicates that the CLIP-SBR framework is highly effective in leveraging and designing inter-session information, leading to better overall performance.
@@
We also applied CLIP-SBR to the inter-session models to assess its effectiveness. The impact varied across the three models: it had a substantial effect on GCSAN, a slight effect on GCEGNN, and almost no effect on LESSR. This suggests that the CLIP-SBR is less beneficial for GCEGNN that already capture inter-session information effectively, while providing significant improvements for GCSAN that do not sufficiently address it. For LESSR, the lack of improvement indicates that the model may not effectively utilize the additional information provided by CLIP-SBR. The impact of CLIP-SBR on inter-session models is further discussed in Section 4.2.3.
@@ 4.2.2. Impact of Item Relationship Mining (RQ2)
Table 3. Comparison of MRR@5 and average training time (in seconds) for the proposed methods and inter-session baselines on the three datasets. Our proposed models are highlighted in gray.
@@ 4.2.3. Impact of Item Cluster-aware Prompt Learning (RQ3)
To evaluate the impact of the Item Cluster-aware Prompt Learning module, we assessed its effectiveness on intra-session and inter-session SBR models. As shown in Table 2 and Figure 2, our approach demonstrates substantial performance improvements, highlighting both the effectiveness and efficiency of the module. In terms of effectiveness, applying the Item Cluster-aware Prompt Learning module led to significant performance improvements in inter-session baselines. Specifically, CLIP-GCSAN showed notable gains, effectively compensating for GCSAN’s limitations in capturing inter-session information, particularly on the Last.fm dataset. For GCEGNN, the module provided moderate improvements, suggesting that while GCEGNN already captures inter-session information effectively, the prompts still offer additional value. Conversely, CLIP-LESSR showed no improvement over the original LESSR, indicating that LESSR’s inherent handling of inter-session information leaves little room for further enhancement from the prompts. Regarding efficiency, the Item Cluster-aware Prompt Learning module demonstrated significant advantages in terms of training time. As shown in Table 3, several CLIP models achieved better or comparable performance (MRR@5) with significantly less training time compared to the inter-session baselines (GCSAN, GCEGNN, LESSR). Among the CLIP models, CLIP-SRGNN shows a good balance between effectiveness and efficiency. It outperformed the inter-session baselines on Last.fm and Reddit, achieving higher MRR@5 with reduced training times—619.79 seconds on Last.fm and 196.71 seconds on Reddit, compared to GCSAN’s 1032.88 seconds, GCEGNN’s 625.71 seconds, and LESSR’s 758.59 seconds on Last.fm, as well as the significantly longer times of the inter-session baselines on Reddit. For the Xing dataset, CLIP-SRGNN achieved comparable performance to the inter-session models while still requiring considerably less time—56.33 seconds versus 92.70 seconds for GCSAN and 60.60 seconds for GCEGNN. The CLIP-TAGNN achieves the best performance on MRR@5 at the expense of longer training time. This highlights the efficiency of the Item Cluster-aware Prompt Learning module, which achieves comparable or superior performance with reduced training time.
@@
4.2.4. Impact of hyperparameter
The Item Relationship Mining module in CLIP-SBR uses the Leiden algorithm for item cluster detection, with the resolution parameter controlling the granularity of community detection. We evaluated CLIP-GRU4Rec using various resolution values ${0.1,0.5,1,1.5,2,3,5}$, with the MRR@5 results shown in Figure 3. Although no clear pattern emerged regarding how resolution changes affected performance, the model consistently performed best with a resolution of 1 across all datasets. Based on these findings, we use a resolution value of 1 in our implementation, as it effectively balances fine-grained and broader item relationships.
@@ 5. Conclusion
In this paper, we introduced the CLuster-aware Item Prompt learning framework for Session-Based Recommendation (CLIP-SBR), aimed at overcoming the limitations of existing SBR methods by effectively capturing both intra- and inter-session item relationships. The framework comprises two key modules: the first module identifies item clusters from session data, while the second module incorporates cluster-specific prompts to enhance the learning capabilities of SBR models. Extensive experiments demonstrated that CLIP-SBR consistently outperforms baseline models, highlighting its effectiveness.

### v2

@@ 4.1.3. Evaluation Metrics
To evaluate the performance of the recommendation, we use two metrics widely adopted in SBR evaluation: Mean Reciprocal Rank (MRR@k) and Recall (Recall@k) following (Ruocco et al., 2017; Li et al., 2017; Quadrana et al., 2017), where $k=5,10$.
@@ 4.1.4. Implementation Details
We implement CLIP-SBR using a popular open-source recommender systems library RecBole [^4] https://recbole.io. All methods are optimized with the Adam optimizer, and we employ early stopping with a patience of 50 epochs to prevent overfitting, using MRR@5 as the indicator. The batch size is tuned from {64, 128, 256, 512, 1024} and the learning rate from {0.001, 0.005, 0.01, 0.05}. For the Leiden algorithm used in item cluster detection, the resolution value is set to 1, with its impact discussed in detail in Section 4.2.4. To ensure robustness, we run each model five times with different random seeds and report the average performance across these runs.
@@ 4.2.1. Overall Performance (RQ1)
Table 2 presents the experimental results of the eight baselines and their corresponding implementations using the CLIP-SBR framework on three real-world datasets. The best result in each column is highlighted in bold. We applied the CLIP-SBR framework to all baselines, including those that focus solely on intra-session relationships and those that incorporate enhancements to better model complex item relationships to assess its impact. The results show that the CLIP-SBR leads to performance improvements across most of the baselines.
@@
In the intra-session baselines, which focus solely on modeling item relationships within indicidual sessions, showed substantial performance improvements with the application of CLIP-SBR, with the most significant gains observed in the NARM model. Among the inter-session baselines, which incorporate mechanisms to better capture item relationships across sessions, CLIP-SBR demonstrated notable improvements in GCSAN, while having a smaller impact on GCEGNN Further analysis of the impact of applying CLIP-SBR to inter-session baselines is provided in Section 4.2.3.
@@
Moreover, the inter-session baselines (GCSAN, GCEGNN) demonstrate generally better performance than the intra-session baselines. Notably, GCEGNN outperforms most intra-session models on the Last.fm dataset. This results highlight the potential of incorporating inter-session information to improve recommendation accuracy, emphasizing the importance of considering both intra-session and inter-session item relationships when developing robust session-based recommendation models.
@@
The application of the CLIP-SBR framework to the intra-session baselines resulted in performance improvements in all case. Specifically, NARM saw a substantial improvement of 45.30% on Last.fm, 20.11% on Xing, and 13.04% on Reddit, marking it as one of the most improved models. TAGNN also demonstrated significant gains, particularly on Last.fm with a 33.91% increase, 15.31% on Xing, and 2.33% on Reddit. Similarly, GRU4Rec achieved notable enhancements of 29.77% on Last.fm, 7.61% on Xing, and 11.10% on Reddit. CORE exhibited improvements of 1.92%, 16.65%, and 4.51% across the respective datasets, while SRGNN showed increases of 31.54%, 1.44%, and 4.26% on the three datasets. SCL showed improvement of 2.33% on Last.fm, 5.18% on Xing, and 0.88% on Reddit. When comparing intra-session CLIP-SBR models with inter-session models, the CLIP-SBR models generally outperformed in many instances. Notably, CLIP-TAGNN consistently surpassed all inter-session models across. CLIP-SRGNN also outperformed inter-session models on Last.fm and Reddit dataset, and outperformed GCSAN in MRR@5 and Recall@5, and GCEGNN in MRR@5. This indicates that the CLIP-SBR framework is highly effective in leveraging and designing inter-session information, leading to better overall performance.
@@
We also applied CLIP-SBR to the inter-session models to assess its effectiveness. The impact varied across the models: it had a substantial effect on GCSAN, and a slight effect on GCEGNN. This suggests that the CLIP-SBR is less beneficial for GCEGNN that already capture inter-session information effectively, while providing significant improvements for GCSAN that do not sufficiently address it. The impact of CLIP-SBR on inter-session models is further discussed in Section 4.2.3.
@@ 4.2.2. Impact of Item Relationship Mining (RQ2)
Table 3. Comparison of MRR@5 and average training time (in seconds) for the baselines and proposed methods on the three datasets. Our proposed models are highlighted in gray.
@@ 4.2.3. Impact of Item Cluster-aware Prompt Learning (RQ3)
To evaluate the impact of the Item Cluster-aware Prompt Learning module, we assessed its effectiveness on intra-session and inter-session SBR models. As shown in Table 2 and Figure 2, our approach demonstrates substantial performance improvements, highlighting both the effectiveness and efficiency of the module. In terms of effectiveness, applying the Item Cluster-aware Prompt Learning module led to significant performance improvements in inter-session baselines. Specifically, CLIP-GCSAN showed notable gains, effectively compensating for GCSAN’s limitations in capturing inter-session information, particularly on the Last.fm dataset. For GCEGNN, the module provided moderate improvements, suggesting that while GCEGNN already captures inter-session information effectively, the prompts still offer additional value.
Regarding efficiency, the Item Cluster-aware Prompt Learning module demonstrated significant advantages in terms of training time. As shown in Table 3, several CLIP models achieved better or comparable performance (MRR@5) with significantly less training time compared to the inter-session baselines (GCSAN, GCEGNN). Note that the reported time in Table 3 accounts only for the Item Cluster-aware Prompt Learning stage, which is the dominant contributor to training cost; the Item Relationship Mining module incurs negligible overhead, taking only about 6 seconds on average.
Among the CLIP models, CLIP-SRGNN shows a good balance between effectiveness and efficiency. It outperformed the inter-session baselines on Last.fm and Reddit, achieving higher MRR@5 with reduced training times—619.79 seconds on Last.fm and 196.71 seconds on Reddit, compared to GCSAN’s 1032.88 seconds, and GCEGNN’s 625.71 seconds on Last.fm, as well as the significantly longer times of the inter-session baselines on Reddit. For the Xing dataset, CLIP-SRGNN achieved comparable performance to the inter-session models while still requiring considerably less time—56.33 seconds versus 92.70 seconds for GCSAN and 60.60 seconds for GCEGNN. The CLIP-TAGNN achieves the best performance on MRR@5 at the expense of longer training time.
This highlights the efficiency of the Item Cluster-aware Prompt Learning module, which achieves comparable or superior performance with reduced training time.
@@
4.2.4. Impact of hyperparameter (RQ4)
The Item Relationship Mining module in CLIP-SBR utilizes the Leiden algorithm for item cluster detection, which includes a key hyperparameter called resolution. The resolution parameter controls the granularity of community detection, thereby affecting the size and number of detected communities. To explore how the resolution value influences the performance of CLIP-SBR, we evaluated the performance of CLIP-GRU4Rec with various resolution values: ${0.1,0.5,1,1.5,2,3,5}$. The results, shown in Figure 3, present the MRR@5 scores across the three datasets. We can observe that while there is no explicit pattern regarding how changes in the resolution value impact performance, the model consistently achieved the best performance when the resolution was set to 1 across all three datasets. This suggests that a resolution value of 1 provides an optimal balance in the granularity of community detection, allowing the model to effectively capture both fine-grained and broader item relationships.
@@ 5. Conclusion
In this paper, we introduced the CLuster-aware Item Prompt learning framework for Session-Based Recommendation (CLIP-SBR), aimed at overcoming the limitations of existing SBR methods by effectively capturing both intra- and inter-session item relationships. The framework comprises two key modules: the first module identifies item clusters from session data, while the second module incorporates cluster-specific prompts to enhance the learning capabilities of SBR models. Extensive experiments demonstrated that CLIP-SBR consistently outperforms baseline models, highlighting its effectiveness. Further analysis of the modules confirmed their significant contributions to the overall performance of CLIP-SBR, with improvements in both recommendation accuracy and computational efficiency. In the future work, we will explore the dynamic prompt learning method that can detect item clusters after each training epoch and assign cluster prompts to each cluster.
GenAI Usage Disclosure
During the preparation of this work, generative AI tools were used to support writing-related tasks such as refining sentence structure, improving clarity and coherence, and assisting with LaTeX formatting for tables and equations. All technical contributions, ideas, experiments, and analyses were solely developed and verified by the authors without any assistance from generative AI systems. No generative AI tools were used for coding, data processing, or result generation. All outputs produced with generative AI assistance were carefully reviewed and edited by the authors to ensure accuracy and adherence to academic standards. This use complies with the ACM policy on authorship and the use of generative AI tools in scholarly publishing.